# A Comprehensive Survey of Speculative Decoding

### v2

@@ 1 Introduction
Speculative Decoding is founded upon two key observations about LLM inference: 1) many easy tokens can be predicted with less computational overhead (e.g., using a smaller model), and 2) LLM inference is highly memory bandwidth bound Patterson (2004) with the main latency bottleneck arising from memory reads/writes of LLM parameters rather than arithmetic computations. Drawing on these observations, Speculative Decoding adapts the concept of speculative execution [^3] Speculative execution  Burton (1985); Hennessy and Patterson (2012) is an optimization technique used in computer architecture where tasks are performed in advance and subsequently verified for their necessity, thereby circumventing the delays inherent in sequential task execution. to focus LLMs’ computational efforts on the validation of pre-drafted tokens, substantially diminishing the need for frequent memory operations of LLM parameters, thereby improving inference efficiency.
While Speculative Decoding shows promise, it raises several critical questions that warrant further investigation. For instance, how to design an optimal drafter to strike a balance between speculation accuracy and drafting efficiency Xia et al. (2022); Li et al. (2024). Additionally, it is essential to assess whether the verification criterion can maintain both generation parallelism and output quality Miao et al. (2023); Cai et al. (2024). Furthermore, since existing Speculative Decoding methods are evaluated under disparate testing conditions, a unified benchmark is needed to facilitate realistic speedup expectations within the research community.
@@ 2 Overview
This paper offers a comprehensive survey of Speculative Decoding. We commence by introducing the early stages of Speculative Decoding research (§3), illustrated by a timeline of its evolution (as shown in Figure 2). This is followed by a formal definition and formulation of Speculative Decoding (§4). Then, we delve into a detailed discussion of leading techniques, including the selection of drafter models (§5), verification strategies (§6), and alignment between the drafter and the target LLM (§7). Moreover, we introduce Spec-Bench, an extensive evaluation benchmark designed for assessing the acceleration effect of Speculative Decoding (§8).
@@ 3.2 Pioneering Draft-then-Verify Efforts
Following SpecDec, Leviathan et al. (2023) and Chen et al. (2023a) made concurrent contributions by proposing Speculative Sampling, expanding this paradigm to encompass the lossless acceleration of nucleus sampling. These methods employed smaller LMs from the same series (e.g., T5-small) to speed up the inference of their larger counterparts (e.g., T5-XXL). Unlike previous work, these off-the-shelf small LMs do not require additional training, enabling the rapid adoption of Speculative Decoding in LLM acceleration. This advancement has elevated Speculative Decoding to the forefront of LLM efficiency research, attracting widespread interest within the NLP community.
@@ 4 Formulation and Definition
1:Target language model $\mathcal{M}_{q}$, drafter model $\mathcal{M}_{p}$, input sequence $x_{1},\dots,x_{t}$, block size $K$, target sequence length $T$, drafting strategy Draft, verification criterion Verify, and correction strategy Correct;
@@
Self-Speculative Zhang et al. (2023a), Lookahead Decoding Fu et al. (2024) , leaf, text width=35em ] ] [ Approximate [ Blockwise Stern et al. (2018), SpecDec Xia et al. (2022), BiLD Kim et al. (2023) , leaf, text width=35em ] ] ] [ Nucleus
@@ 4.1 Autoregressive Decoding
$x_{t+1}\sim q_{n+1}=\mathcal{M}_{q}\left(x\mid x_{<t+1}\right),$ (1)
where $q$ is the conditional probability distribution calculated by $\mathcal{M}_{q}$ and $x_{t+1}$ denotes the next token sampled from $q_{n+1}$. We illustrate a detailed process in Algorithm 1.
@@ 4.2 Speculative Decoding
At each decoding step, Speculative Decoding first efficiently drafts multiple future tokens, as a speculation of the target LLM’s output. Formally, given an input sequence $x_{1},\dots,x_{t}$ and the target LLM $\mathcal{M}_{q}$, this paradigm employs an efficient drafter model $\mathcal{M}_{p}$ (e.g., a smaller LM) to decode the next $K$ drafted tokens:
@@
Methods
$\textsc{Draft}\left(x_{\leq t},\mathcal{M}_{p}\right)$
Drafter Type
Parallel Drafting
$p_{1},\dots,p_{K}=\mathcal{M}_{p}\left(x\mid x_{\leq t}\right)$
FFN Heads Stern et al. (2018); Cai et al. (2024), Non-Autoregressive LM Xia et al. (2022), Mask-Predict Santilli et al. (2023); Fu et al. (2024)
Autoregressive Drafting
$p_{i}=\mathcal{M}_{p}\left(x\mid x_{\leq t},\widetilde{x}_{<i}\right),i=1,\dots,K$
Small LMs Leviathan et al. (2023); Chen et al. (2023a), Early Exiting Yang et al. (2023b), Layer Skipping Zhang et al. (2023a)
@@ 5.1 Independent Drafting
To strike a balance between speculation accuracy and efficiency, SpecDec Xia et al. (2022) first proposed utilizing an independent model for drafting. Specifically, it employed a specialized Non-Autoregressive Transformer that drafts multiple tokens simultaneously per step. This model has a deep-shallow encoder-decoder architecture to run efficiently. Despite its strengths, SpecDec requires training a drafter model from scratch, which demands an increased computational budget.
@@
Methods
$\textsc{Verify}\left(\widetilde{x}_{i},p_{i},q_{i}\right)$
$\textsc{Correct}\left(p_{c},q_{c}\right)$
Representative Work
Greedy Decoding
$\widetilde{x}_{i}=\arg\max q_{i}$
${x}_{t+c}\leftarrow\arg\max q_{c}$
Blockwise Decoding Stern et al. (2018), SpecDec Xia et al. (2022)
Nucleus Sampling
$r<\min\left(1,\frac{q_{i}(\widetilde{x}_{i})}{p_{i}(\widetilde{x}_{i})}\right),r\sim U\left[0,1\right]$
$x_{t+c}\sim\operatorname{norm}(\max\left(0,q_{c}-p_{c}\right))$
Speculative Decoding Leviathan et al. (2023), SpS Chen et al. (2023a)
@@ 5.2 Self-Drafting
While leveraging an external drafter model offers considerable advantages, this approach necessitates extra effort to either train or identify a drafter model that closely aligns with the target LLM. This challenge is intensified when no smaller counterparts of the LLM are available, e.g., LLaMA-7B Touvron et al. (2023a, b). Furthermore, integrating two distinct models within a single system introduces additional computational complexity, particularly in distributed settings Cai et al. (2024).
To address the above issues, numerous studies have suggested leveraging the target LLM itself for efficient drafting Stern et al. (2018); Santilli et al. (2023); Hooper et al. (2023); Cai et al. (2024); Fu et al. (2024); Monea et al. (2023). Particularly, Blockwise Decoding Stern et al. (2018) and Medusa Cai et al. (2024) incorporated FFN heads atop the Transformer decoder, enabling the parallel token generation per step. Compared with external drafters, these lightweight heads reduce extra computational overhead and are friendly to distributed inference. Another line of research has explored the potential of early exiting and layer skipping within the target LLM for drafting Yang et al. (2023b); Zhang et al. (2023a); Hooper et al. (2023). For instance, Yang et al. (2023b) introduced additional subprocesses that exit early during the current decoding step, thereby initiating the drafting of future tokens in advance. Similarly, Self-Speculative Zhang et al. (2023a) proposed to adaptively skip several intermediate layers during inference to draft efficiently.
@@
In each decoding step, the drafted tokens are verified in parallel to ensure the outputs align with the target LLM. This process also determines the number of tokens accepted per step, a vital factor impacting the speedup. This section summarizes various verification criteria $\textsc{Verify}\left(\widetilde{x}_{i},p_{i},q_{i}\right)$ (as shown in Table 2), encompassing those supporting greedy decoding (§6.1) and nucleus sampling (§6.2) in LLM inference. Besides, we introduce token tree verification (§6.3), an effective strategy to increase the token acceptance rate.
@@
Approach Alignment Tuning-free Greedy Nucleus Token Tree
@@
Table 3: Summary of various Speculative Decoding methods. “Independent-D” and “Self-D” denote independent drafting and self-drafting, respectively. “Greedy”, “Nucleus”, and “Token Tree” denote whether the method supports greedy decoding, nucleus sampling, and token tree verification, respectively. We list the most representative target LLMs for each method and the speedups in the original paper (if reported), which is obtained with a batch size of 1.
@@ 6.1 Greedy Decoding
6.2 Nucleus Sampling
Following Stern et al. (2019), subsequent work extended Speculative Decoding to support nucleus sampling Leviathan et al. (2023); Chen et al. (2023a), accelerating the target LLM’s inference without changing its output distribution. Formally, given the initial sequence $x_{1},\dots,x_{t}$, the drafted tokens $\widetilde{x}_{1},\dots,\widetilde{x}_{K}$ and the computed distributions $p_{1},\dots,p_{K}$, $q_{1},\dots,q_{K}$, the verification criterion on the $i_{th}$ drafted token is
@@ 7 Alignment
Though Seq-KD is effective, it ignores the probability distributions of the target LLM, leading to performance degradation when adopted for nucleus sampling. To rectify this, recent studies have explored other KD strategies for Speculative Decoding Zhou et al. (2023); Liu et al. (2023). Notably, DistillSpec Zhou et al. (2023) conducted a comprehensive comparison of different KD strategies on Speculative Decoding across various downstream tasks. Liu et al. (2023) proposed an online KD strategy that dynamically aligns the drafter with the target LLM on the fly using the query data.
@@ 8 Spec-Bench
Figure 5: Speedup comparison of various Speculative Decoding methods on Spec-Bench with greedy settings ($T=0$). Evaluations were conducted on Vicuna-7B with a batch size of 1. We present the mean speedup over 3 runs. The detailed results are shown in Appendix D.
@@ 8.1 Benchmark Construction
To assess Speculative Decoding methods across various scenarios, Spec-Bench encompasses six distinct subtasks: multi-turn conversation, translation, summarization, question answering, mathematical reasoning, and retrieval-augmented generation. We composed Spec-Bench by randomly selecting 80 instances from each of six widely used datasets, including MT-bench Zheng et al. (2023), WMT14 DE-EN, CNN/Daily Mail Nallapati et al. (2016), Natural Questions Kwiatkowski et al. (2019), GSM8K Cobbe et al. (2021), and DPR Karpukhin et al. (2020). For details on Spec-Bench and the specific experimental setup, please refer to Appendix C.
@@ 8.2 Comparative Evaluation
Our main evaluations were conducted on Vicuna-7B at FP16 precision, using a single consumer-grade 3090 GPU [^4] For comparative analysis on a more powerful A100 GPU, please refer to Appendix E.. As depicted in Figure 5, under greedy settings, EAGLE Li et al. (2024) achieves the largest speedup ratio (1.8$\times$×$\sim$2.4$\times$) over autoregressive decoding across most subtasks, especially the mathematical reasoning subtask (with $\sim$2.4$\times$ speedup). PLD Saxena (2023) excels in subtasks with high similarities between input and output, such as summarization (with $\sim$2.4$\times$ speedup). However, its performance diminishes in other subtasks like translation and question answering, with speedup ratios falling between 1.1$\times$×$\sim$1.3$\times$.
@@ 9 Conclusion
This paper presents a comprehensive survey of Speculative Decoding, including a formal definition and formulation, a systematic categorization of existing methods, and an in-depth review of leading techniques. Moreover, we introduce Spec-Bench, an extensive evaluation benchmark for Speculative Decoding methods, and present a comparative evaluation of prominent methods. To our knowledge, this is the first survey dedicated to Speculative Decoding. Our aim for this paper is to clarify the current research landscape and provide insights into future research directions.
@@ Appendix A Applications
Appendix B Challenges and Future Directions
How to trade off speculation accuracy and drafting efficiency?
How to apply Speculative Decoding in batched inference scenarios?
As a general decoding paradigm, Speculative Decoding has already demonstrated its potential in conjunction with other advanced techniques Yang et al. (2023a); Zhang et al. (2023b); Li et al. (2023). For instance, Yuan et al. (2023) combined Speculative Decoding with Contrastive Decoding Li et al. (2023), which not only speeds up the inference but also substantially improves the generation quality. In addition to the acceleration of text-only LLMs, the application of Speculative Decoding in multimodal inference, such as image synthesis, text-to-speech synthesis, and video generation, is also an intriguing and valuable direction for future research. Another promising research direction is to integrate Speculative Decoding with other efficient methods such as vLLM Kwon et al. (2023) and Flash-Attention Dao et al. (2022); Dao (2023), further boosting the inference efficiency of LLM services.
Appendix C Experimental Details
C.1 Details of Spec-Bench

### v3

@@ 1 Introduction
Speculative Decoding is founded upon two key observations about LLM inference: 1) many easy tokens can be predicted with less computational overhead (e.g., using a smaller model), and 2) LLM inference is highly memory bandwidth bound Patterson (2004); Shazeer (2019) with the main latency bottleneck arising from memory reads/writes of LLM parameters rather than arithmetic computations. Drawing on these observations, Speculative Decoding adapts the concept of speculative execution [^3] Speculative execution  Burton (1985); Hennessy and Patterson (2012) is an optimization technique used in computer architecture where tasks are performed in advance and subsequently verified for their necessity, thereby circumventing the delays inherent in sequential task execution. to focus LLMs’ efforts on the validation of pre-drafted tokens, substantially diminishing the need for frequent memory operations of LLM parameters, thereby improving inference efficiency.
While Speculative Decoding shows promise, it raises several critical questions that warrant further investigation. For instance, how to design an optimal drafter to strike a balance between speculation accuracy and drafting efficiency Xia et al. (2023); Zhou et al. (2023); Li et al. (2024). Additionally, it is essential to assess whether the verification criterion can maintain both generation parallelism and output quality Miao et al. (2024); Cai et al. (2024). Furthermore, since existing methods are evaluated under disparate testing conditions, a unified benchmark is needed to facilitate realistic speedup expectations within the research community.
@@ 2 Overview
This paper offers a comprehensive survey of Speculative Decoding. We commence by introducing the early stages of Speculative Decoding research (§3), illustrated by a timeline of its evolution (as shown in Figure 2). This is followed by a formal definition and formulation of Speculative Decoding (§4). Then, we delve into a detailed discussion of leading techniques, including the selection of draft models (§5), verification strategies (§6), and alignment between the drafter and the target LLM (§7). Moreover, we introduce Spec-Bench, an extensive evaluation benchmark designed for assessing the acceleration effect of Speculative Decoding (§8).
@@ 3.2 Pioneering Draft-then-Verify Efforts
Following SpecDec, Leviathan et al. (2023) and Chen et al. (2023a) made concurrent contributions by proposing Speculative Sampling, expanding this paradigm to encompass the lossless acceleration of various sampling methods. These approaches employed smaller LMs from the same series (e.g., T5-small) to speed up the inference of their larger counterparts (e.g., T5-XXL). Unlike previous work, these off-the-shelf small LMs do not require additional training, enabling the rapid adoption of Speculative Decoding in LLM acceleration. This advancement has elevated Speculative Decoding to the forefront of LLM efficiency research, attracting widespread interest within the NLP community.
@@ 4 Formulation and Definition
1:Target language model $\mathcal{M}_{q}$, draft model $\mathcal{M}_{p}$, input sequence $x_{1},\dots,x_{t}$, block size $K$, target sequence length $T$, drafting strategy Draft, verification criterion Verify, and correction strategy Correct;
@@
Self-Speculative Zhang et al. (2023a), Lookahead Decoding Fu et al. (2024) , leaf, text width=35em ] ] [ Approximate [ Blockwise Stern et al. (2018), SpecDec Xia et al. (2023), BiLD Kim et al. (2023) , leaf, text width=35em ] ] ] [ Speculative
@@ 4.1 Autoregressive Decoding
$x_{t+1}\sim q_{t+1}=\mathcal{M}_{q}\left(x\mid x_{<t+1}\right),$ (1)
where $q$ is the conditional probability distribution calculated by $\mathcal{M}_{q}$ and $x_{t+1}$ denotes the next token sampled from $q_{t+1}$. We illustrate a detailed process in Algorithm 1.
@@ 4.2 Speculative Decoding
At each decoding step, Speculative Decoding first efficiently drafts multiple future tokens, as a speculation of the target LLM’s output. Formally, given an input sequence $x_{1},\dots,x_{t}$ and the target LLM $\mathcal{M}_{q}$, this paradigm employs an efficient draft model $\mathcal{M}_{p}$ (e.g., a smaller LM) to decode the next $K$ drafted tokens:
@@
Methods $\textsc{Draft}\left(x_{\leq t},\mathcal{M}_{p}\right)$ Drafter Type
Parallel Drafting $p_{1},\dots,p_{K}=\mathcal{M}_{p}\left(x\mid x_{\leq t}\right)$ FFN Heads Stern et al. (2018); Cai et al. (2024), Non-Autoregressive LM Xia et al. (2023), Mask-Predict Santilli et al. (2023); Fu et al. (2024)
Autoregressive Drafting $p_{i}=\mathcal{M}_{p}\left(x\mid x_{\leq t},\widetilde{x}_{<i}\right),i=1,\dots,K$ Small LMs Leviathan et al. (2023); Chen et al. (2023a), Early Exiting Yang et al. (2023b), Layer Skipping Zhang et al. (2023a)
@@ 5.1 Independent Drafting
To strike a balance between speculation accuracy and efficiency, SpecDec Xia et al. (2023) first proposed utilizing an independent model for drafting. Specifically, it employed a specialized Non-Autoregressive Transformer that drafts multiple tokens simultaneously per step. This model has a deep-shallow encoder-decoder architecture to run efficiently. Despite its strengths, SpecDec requires training a draft model from scratch, which demands an increased computational budget.
@@
Methods $\textsc{Verify}\left(\widetilde{x}_{i},p_{i},q_{i}\right)$ $\textsc{Correct}\left(p_{c},q_{c}\right)$ Representative Work
Greedy Decoding $\widetilde{x}_{i}=\arg\max q_{i}$ ${x}_{t+c}\leftarrow\arg\max q_{c}$ Blockwise Decoding Stern et al. (2018), SpecDec Xia et al. (2023)
Speculative Sampling $r<\min\left(1,\frac{q_{i}(\widetilde{x}_{i})}{p_{i}(\widetilde{x}_{i})}\right),r\sim U\left[0,1\right]$ $x_{t+c}\sim\operatorname{norm}(\max\left(0,q_{c}-p_{c}\right))$ Speculative Decoding Leviathan et al. (2023), SpS Chen et al. (2023a)
@@ 5.2 Self-Drafting
While leveraging an external draft model offers considerable advantages, this approach necessitates extra effort to either train or identify a draft model that closely aligns with the target LLM. This challenge is intensified when no smaller counterparts of the LLM are available, e.g., LLaMA-7B Touvron et al. (2023a, b). Furthermore, integrating two distinct models within a single system introduces additional computational complexity, particularly in distributed settings Cai et al. (2024).
To address the above issues, numerous studies have suggested leveraging the target LLM itself for efficient drafting Stern et al. (2018); Santilli et al. (2023); Hooper et al. (2023); Cai et al. (2024); Fu et al. (2024); Du et al. (2024). Particularly, Blockwise Decoding Stern et al. (2018) and Medusa Cai et al. (2024) incorporated FFN heads atop the Transformer decoder, enabling the parallel token generation per step. Compared with external drafters, these lightweight heads reduce extra computational overhead and are friendly to distributed inference. Another line of research has explored the potential of early exiting and layer skipping within the target LLM for drafting Yang et al. (2023b); Zhang et al. (2023a); Hooper et al. (2023). For instance, Yang et al. (2023b) introduced additional subprocesses that exit early during the current decoding step, thereby initiating the drafting of future tokens in advance. Similarly, Self-Speculative Zhang et al. (2023a) proposed to adaptively skip several intermediate layers during inference to draft efficiently.
@@
In each decoding step, the drafted tokens are verified in parallel to ensure the outputs align with the target LLM. This process also determines the number of tokens accepted per step, a vital factor impacting the speedup. This section summarizes various verification criteria $\textsc{Verify}\left(\widetilde{x}_{i},p_{i},q_{i}\right)$ (as shown in Table 2), encompassing those supporting greedy decoding (§6.1) and speculative sampling (§6.2) in LLM inference. Besides, we introduce token tree verification (§6.3), an effective strategy to increase the token acceptance rate.
@@
Approach Alignment Tuning-free Greedy Sampling Token Tree
@@
Table 3: Summary of Speculative Decoding methods. “Independent-D” and “Self-D” denote independent drafting and self-drafting, respectively. “Greedy”, “Sampling”, and “Token Tree” denote whether the method supports greedy decoding, speculative sampling, and token tree verification, respectively. We list the most representative target LLMs for each method and the speedups in the original paper (if reported), which is obtained with a batch size of 1.
@@ 6.1 Greedy Decoding
6.2 Speculative Sampling
Following Stern et al. (2019), subsequent work extended Speculative Decoding to support various sampling methods Leviathan et al. (2023); Chen et al. (2023a), accelerating the target LLM’s inference without changing its output distribution. Formally, given the initial sequence $x_{1},\dots,x_{t}$, the drafted tokens $\widetilde{x}_{1},\dots,\widetilde{x}_{K}$ and the computed distributions $p_{1},\dots,p_{K}$, $q_{1},\dots,q_{K}$, the verification criterion on the $i_{th}$ drafted token is
@@ 7 Alignment
Although Seq-KD is effective, it ignores the probability distributions of the target LLM, leading to performance degradation with sampling methods. To rectify this, recent studies have explored other KD strategies for Speculative Decoding Zhou et al. (2023); Liu et al. (2023). Notably, DistillSpec Zhou et al. (2023) conducted a comprehensive comparison of different KD strategies on Speculative Decoding across various downstream tasks. Liu et al. (2023) proposed an online KD strategy that dynamically aligns the drafter with the target LLM on the fly using the query data.
@@ 8.1 Benchmark Construction
To assess Speculative Decoding methods across various scenarios, Spec-Bench encompasses six distinct subtasks: multi-turn conversation, translation, summarization, question answering, mathematical reasoning, and retrieval-augmented generation. We composed Spec-Bench by randomly selecting 80 instances from each of six widely used datasets, including MT-bench Zheng et al. (2023), WMT14 DE-EN, CNN/Daily Mail Nallapati et al. (2016), Natural Questions Kwiatkowski et al. (2019), GSM8K Cobbe et al. (2021), and DPR Karpukhin et al. (2020). For details on Spec-Bench and the specific experimental setup, please refer to Appendix B.
Figure 5: Speedup comparison of various Speculative Decoding methods on Spec-Bench with greedy settings ($T=0$). Evaluations were conducted on Vicuna-7B with a batch size of 1. We present the mean speedup over 3 runs. The detailed results are shown in Appendix C.
@@ 8.2 Comparative Evaluation
Our main evaluations were conducted on Vicuna-7B at FP16 precision using a single consumer-grade 3090 GPU [^4] For comparative analysis on a more powerful A100 GPU, please refer to Appendix D.. As depicted in Figure 5, under greedy settings, EAGLE Li et al. (2024) achieves the highest speedup ratio (1.8$\times$×$\sim$2.4$\times$) over autoregressive decoding across most subtasks, especially in mathematical reasoning (with a $\sim$2.4$\times$ speedup). EAGLE’s success is mainly due to two factors: 1) it reuses the KV cache of LLMs to predict drafted tokens, substantially reducing the drafting computational overhead; 2) compared with Medusa Cai et al. (2024), EAGLE drafts in an autoregressive way, providing more stable and accurate speculation results. PLD Saxena (2023) excels in subtasks with high similarities between input and output, such as summarization (with a $\sim$2.4$\times$ speedup). However, its performance diminishes in other subtasks like translation and question answering, with speedup ratios falling between 1.1$\times$×$\sim$1.3$\times$.
@@
9 Challenges and Future Directions
How to trade off speculation accuracy and drafting efficiency?
How to apply Speculative Decoding in batched inference scenarios?
As a general decoding paradigm, Speculative Decoding has already demonstrated its potential in conjunction with other advanced techniques Yang et al. (2023a); Zhang et al. (2023b); Li et al. (2023). For instance, Yuan et al. (2023) combined Speculative Decoding with Contrastive Decoding Li et al. (2023), which not only speeds up the inference but also substantially improves the generation quality. In addition to the acceleration of text-only LLMs, applying Speculative Decoding in multimodal inference, such as image synthesis, text-to-speech synthesis, and video generation, is also an intriguing and valuable direction for future research. Another promising research direction is to integrate Speculative Decoding with other efficient methods such as vLLM Kwon et al. (2023), Non-Auregressive Generation Du et al. (2021, 2022) and Flash-Attention Dao et al. (2022); Dao (2023), further boosting the inference efficiency of LLM services.
@@ 10 Conclusion
This paper presents a comprehensive survey of Speculative Decoding, including the evolution of this promising paradigm, its formal definition and formulation, a systematic categorization of existing methods, and an in-depth review of leading techniques. Moreover, we introduce Spec-Bench, an extensive evaluation benchmark for Speculative Decoding methods, and present a comparative evaluation of prominent methods. To our knowledge, this is the first survey dedicated to Speculative Decoding. Our aim for this paper is to clarify the current research landscape and provide insights into future research directions.
@@
Acknowledgements
We thank all anonymous reviewers for their valuable comments during the review process. This work is partially supported by Research Grants Council of Hong Kong (15207122 and 15213323).
@@ Appendix A Applications
Appendix B Experimental Details
B.1 Details of Spec-Bench